# Two-generation of traceless matrices over finite fields

### v1

@@ Abstract.
We prove that the Lie algebra $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ of traceless matrices over a finite field of odd characteristic $p$ can be generated by $2$ elements with a single exception $(n,p)=(3,3)$. We establish a curious trace identity in $\operatorname{\mathfrak{sl}}_{3}(\mathbf{F})$ over any field $\mathbf{F}$ of characteristic $3$ that obstructs $2$-generation.
@@ 1. Introduction
This $2$-generation phenomenon is well-established across various classical algebraic structures over classical fields such as the complex numbers. For simple associative algebras, this property is underscored by the classical Burnside irreducibility theorem, which posits that over algebraically closed fields, any two matrices without a common proper invariant subspace generate the whole matrix algebra. Similarly, for simple algebraic groups, it is a result of Guralnick [Gur98, Theorem 3.3] that over an algebraically closed field of any characteristic, a simple algebraic group $G$ is topologically $3/2$-generated, meaning that for every nontrivial $x$ there exists $y$ so that $\langle x,y\rangle$ is dense in $G$. Furthermore, simple Lie algebras over any field of characteristic 00 have been shown to be $2$-generated by Kuranishi [Kur51], with extensions to $3/2$-generation in characteristic 00 by Ionescu [Ion76] and in positive characteristic $p>3$ by Bois [Boi09]. Recently, Chistopolskaya [Chi18] has explored nilpotent pairs of generators in traceless matrices.
@@
Let $\mathbf{F}_{q}$ be a finite field of odd characteristic $p$. Then the simple Lie algebra $\operatorname{\mathfrak{psl}}_{n}(\mathbf{F}_{q})$ is $2$-generated as long as $(n,p)\neq(3,3)$.
We show that $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$ is, intriguingly, not $2$-generated when $\mathbf{F}$ is any field of characteristic $3$, although it does allow for $3$-generation. This is, to the best of our knowledge, the first example of a simple Lie algebra over an infinite field that is not generated by two elements.
The strategy of the proof is outlined as follows. Our argument rests on the concept of consistent matrices (Section 2) in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$. We explain what these matrices are, where they come from and how they can be used to prove $2$-generation of $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$, and consequently of $\mathfrak{psl}_{n}(\mathbf{F}_{q})$, as long as they exist. When $n$ is bounded and the field $\mathbf{F}_{q}$ is large enough (this is the case of bounded rank, Section 3), we can produce these matrices from suitable subsets of integers related to distinct sum sets from additive combinatorics. On the other hand, when $n$ tends to infinity and the field is small (this is the case of unbounded rank, Section 4), we produce generating pairs from companion matrices of suitable elements in extensions of $\mathbf{F}_{q}$. In the simple case, when $p$ does not divide $n$, normal elements do the job. In the modular case, when $p$ divides $n$, we replace normal elements by developing their traceless version which we call sharply traceless elements. The single exceptional case in odd characteristic (Section 5) that our method does not cover is the Lie algebra $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F}_{q})$ in characteristic $3$. We demonstrate that these algebras are not $2$-generated by establishing a curious identity between any two elements of $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F}_{q})$ that obstructs $2$-generation. The same phenomenon occurs in $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F}_{q})$ in characteristic $2$ (Appendix A). Supported by computer experiments, we conjecture this is the only exception in even characteristic.
@@ 2. Consistent matrices
with distinct nonzero diagonal terms over a field $\mathbf{F}$ satisfying the condition
$\lambda_{i}-\lambda_{j}=\lambda_{k}-\lambda_{l}\textup{ if and only if }(i,j)=(k,l)\text{ or }(i,k)=(j,l).$
@@
Note that when $\operatorname{char}(\mathbf{F})=2$, one has $\lambda_{i}-\lambda_{j}=\lambda_{j}-\lambda_{i}$ for any $i,j$, so no consistent set exists.
@@ 3. Bounded rank
Using consistent matrices, we can easily find $2$ generators for Lie algebras $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ of bounded rank, meaning that $n\geq 2$ is fixed and either the characteristic $p$ of the field $\mathbf{F}_{q}$ is large enough in terms of $n$ or $q$ is large enough in terms of $p$ and $n$. We assume throughout that $p>2$, since no consistent sets exist in even characteristic.
@@ 3.1. Bounded characteristic and large field
A sharply traceless set of size $n$ is consistent if and only if $(n,p)\neq(3,3)$.
@@
The relation violating consistency $s_{i}-s_{j}=s_{k}-s_{l}$ involves at most $4$ elements, so the set is consistent if $n\geq 5$. When $n=4$, there are no linear relations involving at most $3$ terms, so a violating relation can be a scalar multiple of the sum zero relation only when $p=2$. Similarly, when $n=3$, consistency can be violated only when $p=2$ or $p=3$ (for example in the form $s_{1}-s_{2}=s_{2}-s_{3}$, a relation that is equivalent to sum zero). Conversely, in case $(n,p)=(3,3)$, a sharply traceless set $\{s_{1},s_{2},s_{3}\}$ always satisfies $s_{3}-s_{2}=s_{2}-s_{1}$, so it is not consistent. Finally, when $n=2$, a sharply traceless set $\{s,-s\}$ with $s\neq 0$ is consistent provided $p\neq 2$. ∎
It is not difficult to construct sharply traceless sets as long as the field is large enough. [^1] It is, however, much more challenging to construct sharply traceless sets over small fields. We will do this in the next section.
@@
Let $(n,p)\neq(3,3)$ and $q\geq p^{n-1}$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@ 4.1. Polynomials with consistent roots
The coefficients of $P$ and $P^{-1}\mathbf{1}P$ are related to the field trace of the extension $\mathbf{F}_{q^{n}}$ of $\mathbf{F}_{q}$, which computes, for a given $\alpha\in\mathbf{F}_{q^{n}}$, the sum of its Galois conjugates:
@@
Let $\mathbf{F}$ be a splitting field for $f$. Let $C$ be the companion matrix of $f$. Since the roots of $f$ form a consistent set, we have $C\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ and the eigenvalues of $C$, being the roots of $f$, are all distinct in $\mathbf{F}$. Therefore $C$ diagonalises to a consistent matrix $D=PCP^{-1}$ for some $P\in\operatorname{GL}_{n}(\mathbf{F})$, which generates $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ together with the matrix $\mathbf{1}$. Thus, according to [Boi09, Proposition 1.1.3], there exist $n^{2}-1$ Lie monomials $\{g_{i}(X,Y)\}_{i}$ in two variables whose evaluations in $(D,\mathbf{1})$ span the vector space $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Since Lie polynomials are equivariant, [^2] Meaning that if $g(X,Y)$ is a Lie polynomial, then $P^{-1}g(X,Y)P=g(P^{-1}XP,P^{-1}YP)$. the matrices $g_{i}(C,P^{-1}\mathbf{1}P)=P^{-1}g_{i}(D,\mathbf{1})P$ also span $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Now, both $C$ and $P^{-1}\mathbf{1}P$ belong to $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$, and as each $g_{i}$ is a Lie monomial, its evaluation $g_{i}(C,P^{-1}\mathbf{1}P)$ is in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$. Therefore $C$ and $P^{-1}\mathbf{1}P$ generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ as a Lie algebra over $\mathbf{F}_{q}$. ∎
@@ 4.2. Normal elements
Let $n$ be coprime to $p$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@ 4.3. Sharply traceless elements
In case $p$ divides $n$, normal elements still exist, but there is no evident way of producing a traceless matrix from it.
@@
Let $(n,p)\neq(3,3)$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@
A sharply traceless elements gives a sharply traceless set of its Galois conjugates. The result follows from Lemma 3.1. ∎
Our proof of Theorem 4.5 mimics the usual approach for finding a normal element, compare with [vzGG90] or [LN83] and the closely related notion of a cyclic element from [NP95]. The argument rests on inspecting the Frobenius endomorphism. More precisely, we consider various polynomials in $\operatorname{Fr}$ and give a characterization of sharply traceless elements in $\mathbf{F}_{q^{n}}$ in terms of certain divisors of $m(x)=x^{n}-1$, the minimal and at the same time characteristic polynomial of $\operatorname{Fr}$ (see [LN83, proof of Theorem 2.35]). The most important divisor of $m$ for us is
@@
since, for an element $\alpha\in\mathbf{F}_{q^{n}}$, we have
$\operatorname{Tr}(\alpha)=\alpha^{q^{n-1}}+\alpha^{q^{n-2}}+\dots+\alpha=\Phi(\operatorname{Fr})\alpha.$
The other divisors that will be relevant in describing sharply traceless elements are obtained in the following manner. Let $\operatorname{Div}(\Phi)$ be the set of prime factors of $\Phi$, i.e., distinct monic irreducible polynomials in $\mathbf{F}_{q}[x]$ that divide $\Phi$. For a given $\pi\in\operatorname{Div}(\Phi)$, let $\operatorname{mult}(\pi)$ be the multiplicity of $\pi$ in $\Phi$ and set
$\Phi_{\pi}(x)=\frac{\Phi(x)}{\pi(x)}.$
Proposition 4.7.
An element $\alpha\in\mathbf{F}_{q^{n}}$ is sharply traceless if and only if
$\Phi({\operatorname{Fr}})\alpha=0\quad\text{and}\quad\Phi_{\pi}({\operatorname{Fr}})\alpha\neq 0\text{ for every $\pi\in\operatorname{Div}(\Phi)$.}$
@@
It is clear that a sharply traceless element $\alpha$ satisfies $\Phi(\operatorname{Fr})\alpha=\operatorname{Tr}(\alpha)=0$ and, as there is no linear relation between the Galois conjugates of $\alpha$ having less than $n$ terms, we must also have $\Phi_{\pi}(\operatorname{Fr})\alpha\neq 0$.
As for the converse, assume $\alpha$ is a traceless element with $\Phi_{\pi}(\operatorname{Fr})\alpha\neq 0$ for each $\pi\in\operatorname{Div}(\Phi)$. Suppose now that $\alpha$ satisfies a nontrivial linear dependence relation over $\mathbf{F}_{q}$, i.e.,
$\sum_{i=1}^{n-1}a_{i}\alpha^{q^{i}}=0$
for some $a_{i}\in\mathbf{F}_{q}$. In other words, the polynomial $d(x)=\sum_{i=1}^{n}a_{i}x^{i}\in\mathbf{F}_{q}[x]$ has the property $d({\operatorname{Fr}})\alpha=0$. We can assume $d$ is the monic polynomial of minimal degree with this property. Since both $\Phi(\operatorname{Fr})$ and $d(\operatorname{Fr})$ kill $\alpha$, it follows by Bézout that $\gcd(\Phi,d)(\operatorname{Fr})$ does the same. Minimality of $d$ implies that $d=\gcd(\Phi,d)$, so $d$ divides $\Phi$. On the other hand, $d$ does not divide $\Phi_{\pi}$ as $\Phi_{\pi}(\operatorname{Fr})\alpha\neq 0$. It follows that $d=\Phi$, and so the linear relation above is just $\operatorname{Tr}(\alpha)=0$ up to a scalar, proving that $\alpha$ is sharply traceless. ∎
We now use this criterion for sharply traceless elements by decomposing the vector space $\mathbf{F}_{q^{n}}$ into root subspaces of primes divisors of $\Phi$ evaluated in $\operatorname{Fr}$. For a given $\pi\in\operatorname{Div}(\Phi)$, let
$V_{\pi}=\ker\pi^{\operatorname{mult}(\pi)}({\operatorname{Fr}})\quad\text{and}\quad W_{\pi}=\ker\pi^{\operatorname{mult}(\pi)-1}({\operatorname{Fr}}).$
Let $X_{\pi}$ be a complement to $W_{\pi}$ in $V_{\pi}$. We thus have the decomposition
$\ker\Phi({\operatorname{Fr}})=\bigoplus_{\pi\in\operatorname{Div}(\Phi)}V_{\pi}=\bigoplus_{\pi\in\operatorname{Div}(\Phi)}\left(W_{\pi}\oplus X_{\pi}\right).$
Consider the particular divisor $x-1$, whose multiplicity in $m(x)$ is $\operatorname{mult}(x-1)+1$. Let $U$ be a complement to $V_{x-1}$ in $\ker(x-1)^{\operatorname{mult}(x-1)+1}(\operatorname{Fr})$. Then the whole space $\mathbf{F}_{q^{n}}=\ker m(\operatorname{Fr})$ decomposes as
(2) $\mathbf{F}_{q^{n}}=U\oplus\bigoplus_{\pi\in\operatorname{Div}(\Phi)}\left(W_{\pi}\oplus X_{\pi}\right).$
This decomposition can be used to characterize sharply traceless elements as follows.
Proposition 4.8.
Let $\alpha\in\mathbf{F}_{p^{n}}$ with a decomposition according to (2) as
$\alpha=u+\sum_{\pi\in\operatorname{Div}(\Phi)}\left(w_{\pi}+x_{\pi}\right).$
Then $\alpha$ is sharply traceless if and only if $u=0$ and $x_{\pi}\neq 0$ for every $\pi\in\operatorname{Div}(\Phi)$.
Proof.
@@
Suppose first that $\alpha$ is sharply traceless. Certainly $u=0$, since $\operatorname{Tr}(\alpha)=0$ and so $\alpha\in\ker\Phi(\operatorname{Fr})$. Write $v_{\pi}=w_{\pi}+x_{\pi}\in V_{\pi}$, so that
$\alpha=\sum_{\pi\in\operatorname{Div}(\Phi)}v_{\pi}.$
Note that for any $\pi^{\prime}\neq\pi$, we have $\Phi_{\pi}(\operatorname{Fr})v_{\pi^{\prime}}=0$ since $\Phi_{\pi}$ is divisible by $\pi^{\prime\operatorname{mult}(\pi^{\prime})}$. Hence $\Phi_{\pi}({\operatorname{Fr}})\alpha=\Phi_{\pi}({\operatorname{Fr}})v_{\pi}$. Supposing $x_{\pi}=0$, it follows that $v_{\pi}\in W_{\pi}$, and since $W_{\pi}\leq\ker\Phi_{\pi}(\operatorname{Fr})$, we obtain $\Phi_{\pi}(\operatorname{Fr})\alpha=0$. This is impossible by Proposition 4.7.
Assume now that $u=0$ and $x_{\pi}\neq 0$ for every $\pi$. Write $v_{\pi}=w_{\pi}+x_{\pi}$ as above. In order to prove that $\alpha$ is sharply traceless, we again rely on Proposition 4.7. First of all, since $\pi^{\operatorname{mult}(\pi)}$ divides $\Phi$, we have $\Phi(\operatorname{Fr})v_{\pi}=0$ and so $\Phi(\operatorname{Fr})\alpha=0$. Now consider $\Phi_{\pi}(\operatorname{Fr})\alpha$. As above, we have $\Phi_{\pi}({\operatorname{Fr}})\alpha=\Phi_{\pi}({\operatorname{Fr}})v_{\pi}$. The greatest common divisor of $\Phi_{\pi}$ and $\pi^{\operatorname{mult}(\pi)}$ is $\pi^{\operatorname{mult}(\pi)-1}$, therefore by Bézout there exist polynomials $a,b\in\mathbf{F}_{q}[x]$ with
$\pi^{\operatorname{mult}(\pi)-1}=a\Phi_{\pi}+b\pi^{\operatorname{mult}(\pi)}.$
Evaluating at $\operatorname{Fr}$ and acting on $v_{\pi}$ gives
$\pi^{\operatorname{mult}(\pi)-1}(\operatorname{Fr})v_{\pi}=a\Phi_{\pi}(\operatorname{Fr})v_{\pi}.$
The left hand side is nonzero as $x_{\pi}\neq 0$ and so $v_{\pi}\notin W_{\pi}$, and the right hand side equals $a\Phi_{\pi}(\operatorname{Fr})\alpha$. Hence $\Phi_{\pi}(\operatorname{Fr})\alpha\neq 0$, as required. ∎
Dimensions of the subspaces appearing in decomposition (2) can be computed using the following well-known lemma.
Lemma 4.9 ([vzGG90, Lemma 4.1]).
Let $A$ be an endomorphism of a finite dimensional vector space over a field $\mathbf{F}$ whose minimal polynomial $m$ coincides with the characteristic polynomial. For any divisor $d\in F[x]$ of $m$, we have $\dim\ker d(A)=\deg d$.
The lemma gives that
$\dim W_{\pi}=(\operatorname{mult}(\pi)-1)\cdot\deg(\pi),\quad\dim X_{\pi}=\deg(\pi).$
The number of sharply traceless elements in $\mathbf{F}_{q^{n}}$ is therefore equal to
(3) $\prod_{\pi\in\operatorname{Div}(\Phi)}\left(|W_{\pi}|\cdot|X_{\pi}-\{0\}|\right)=\prod_{\pi\in\operatorname{Div}(\Phi)}q^{(\operatorname{mult}(\pi)-1)\deg\pi}(q^{\deg(\pi)}-1).$
In particular, this number is positive and so sharply traceless elements indeed exist. Furthermore, we have
$\sum_{\pi\in\operatorname{Div}(\Phi)}(\operatorname{mult}(\pi)-1)\deg(\pi)=\deg(\Phi)-\sum_{\pi\in\operatorname{Div}(\Phi)}\deg(\pi)$
and therefore (3) simplifies into the following.

### v2

@@ Abstract.
We prove that the Lie algebra $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ of traceless matrices over a finite field of characteristic $p$ can be generated by $2$ elements with exceptions when $(n,p)$ is $(3,3)$ or $(4,2)$. In the latter cases, we establish curious identities that obstruct $2$-generation.
@@ 1. Introduction
This $2$-generation phenomenon is well-established across various classical algebraic structures over classical fields such as the complex numbers. For simple associative algebras, this property is underscored by the classical Burnside irreducibility theorem, which posits that over algebraically closed fields, any two matrices without a common proper invariant subspace generate the whole matrix algebra. Similarly, for simple algebraic groups, it is a result of Guralnick [Gur98, Theorem 3.3] that over an algebraically closed field of any characteristic, a simple algebraic group $G$ is topologically $3/2$-generated, meaning that for every nontrivial $x$ there exists $y$ so that $\langle x,y\rangle$ is dense in $G$. Furthermore, simple Lie algebras over any field of characteristic 00 have been shown to be $2$-generated by Kuranishi [Kur51] (see also [BO76] for a more generic approach), and explicit generating pairs were found by Grozman and Leites [LG96]. This was extended to $3/2$-generation in characteristic 00 by Ionescu [Ion76] and in positive characteristic $p>3$ by Bois [Boi09]. More recently, Chistopolskaya [Chi18] has explored nilpotent pairs of generators in traceless matrices.
@@
Let $\mathbf{F}_{q}$ be a finite field of characteristic $p$. Then the simple Lie algebra $\operatorname{\mathfrak{psl}}_{n}(\mathbf{F}_{q})$ is $2$-generated as long as $(n,p)\neq(3,3),(4,2)$.
We show that both $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F})$ and $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F})$ are, intriguingly, not $2$-generated when $\mathbf{F}$ is any field of characteristic $3$ and $2$ respectively. However, these Lie algebras do allow for $3$-generation. These are, to the best of our knowledge, the first known examples of simple Lie algebras over infinite fields that are not generated by two elements, thus answering [GKPS06, Question 2.3] as well as [Chi18, Problem 2].
The strategy of the proof is outlined as follows. Our argument rests on the concept of consistent matrices (Section 2) in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$. We explain what these matrices are, where they come from and how they can be used to prove $2$-generation of $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$, and consequently of $\mathfrak{psl}_{n}(\mathbf{F}_{q})$, as long as they exist. When $n$ is fixed and the field $\mathbf{F}_{q}$ is large enough in terms of $n$ (this is the case of bounded rank, Section 3), we can produce these matrices from suitable subsets of integers related to distinct sum sets from additive combinatorics. On the other hand, when $n$ tends to infinity and the field is small (this is the case of unbounded rank, Section 4), we produce generating pairs from companion matrices of suitable elements in extensions of $\mathbf{F}_{q}$. In the simple case, when $p$ does not divide $n$, normal elements do the job. In the modular case, when $p$ divides $n$, we replace normal elements by developing their traceless version which we call sharply traceless elements. Even characteristic (Section 5) requires special care and a further step to reach $2$-generation. The exceptional cases (Section 6) that our method does not cover are the Lie algebras $\operatorname{\mathfrak{psl}}_{3}(\mathbf{F}_{q})$ and $\operatorname{\mathfrak{psl}}_{4}(\mathbf{F}_{q})$ in characteristics $3$ and $2$ respectively. We demonstrate that these algebras are not $2$-generated by establishing a curious identity between any two elements in these Lie algebras that obstructs $2$-generation.
@@ 2. Consistent matrices
over a field $\mathbf{F}$ satisfying the condition
(1) $\lambda_{i}-\lambda_{j}=\lambda_{k}-\lambda_{l}\ \textup{ if and only if }\ (i,j)=(k,l)\text{ or }(i,k)=(j,l).$
@@
Note that when $\operatorname{char}(\mathbf{F})=2$, one has $\lambda_{i}-\lambda_{j}=\lambda_{j}-\lambda_{i}$ for any $i,j$, so no consistent set exists. We will deal with even characteristic separately in Section 5.
@@ 3. Bounded rank
Using consistent matrices, we can easily find $2$ generators for Lie algebras $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ of bounded rank, meaning that $n\geq 2$ is fixed and either the characteristic $p$ of the field $\mathbf{F}_{q}$ is large enough in terms of $n$ or $q$ is large enough in terms of $p$ and $n$.
@@ 3.1. Bounded characteristic and large field
A sharply traceless set of size $n$ is consistent if and only if $p\neq 2$ and $(n,p)\neq(3,3)$.
@@
No consistent sets exist in even characteristic. Assume now that $p\neq 2$. The relation violating consistency $s_{i}-s_{j}=s_{k}-s_{l}$ involves at most $4$ elements, so the set is consistent if $n\geq 5$. When $n=4$, there are no linear relations involving at most $3$ terms, so a violating relation can be a scalar multiple of the sum zero relation only when $p=2$. Similarly, when $n=3$, consistency can be violated only when $p=2$ or $p=3$ (for example in the form $s_{1}-s_{2}=s_{2}-s_{3}$, a relation that is equivalent to sum zero). Conversely, in case $(n,p)=(3,3)$, a sharply traceless set $\{s_{1},s_{2},s_{3}\}$ always satisfies $s_{3}-s_{2}=s_{2}-s_{1}$, so it is not consistent. Finally, when $n=2$, a sharply traceless set $\{s,-s\}$ with $s\neq 0$ is consistent provided $p\neq 2$. ∎
It is not difficult to construct sharply traceless sets as long as the field is large enough.
@@
Let $p\neq 2$, $(n,p)\neq(3,3)$, $q\geq p^{n-1}$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@ 4.1. Polynomials with consistent roots
The entries of $P$ and $P^{-1}\mathbf{1}P$ are related to the field trace of the extension $\mathbf{F}_{q^{n}}$ of $\mathbf{F}_{q}$, which computes, for a given $\alpha\in\mathbf{F}_{q^{n}}$, the sum of its Galois conjugates:
@@
Let $\mathbf{F}$ be a splitting field for $f$. Let $C$ be the companion matrix of $f$. Since the roots of $f$ form a consistent set, we have $C\in\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ and the eigenvalues of $C$, being the roots of $f$, are all distinct in $\mathbf{F}$. Therefore $C$ diagonalizes to a consistent matrix $D=PCP^{-1}$ for some $P\in\operatorname{GL}_{n}(\mathbf{F})$, which generates $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ together with the matrix $\mathbf{1}$. Thus, according to [Boi09, Proposition 1.1.3], there exist $n^{2}-1$ Lie monomials $\{g_{i}(X,Y)\}_{i}$ in two variables whose evaluations in $(D,\mathbf{1})$ span the vector space $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Since Lie polynomials are equivariant, [^1] Meaning that if $g(X,Y)$ is a Lie polynomial, then $P^{-1}g(X,Y)P=g(P^{-1}XP,P^{-1}YP)$. the matrices $g_{i}(C,P^{-1}\mathbf{1}P)=P^{-1}g_{i}(D,\mathbf{1})P$ also span $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F})$ over $\mathbf{F}$. Now, both $C$ and $P^{-1}\mathbf{1}P$ belong to $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$, and as each $g_{i}$ is a Lie monomial, its evaluation $g_{i}(C,P^{-1}\mathbf{1}P)$ is in $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$. Therefore $C$ and $P^{-1}\mathbf{1}P$ generate $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ as a Lie algebra over $\mathbf{F}_{q}$. ∎
@@ 4.2. Normal elements
Let $p\neq 2$ and $n$ coprime to $p$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@ 4.3. Sharply traceless elements
In case $p$ divides $n$, normal elements still exist, but there is no evident way of producing a traceless matrix from them.
@@
Let $p\neq 2$ and $(n,p)\neq(3,3)$. Then $\operatorname{\mathfrak{sl}}_{n}(\mathbf{F}_{q})$ is $2$-generated.
@@
A sharply traceless element gives a sharply traceless set of its Galois conjugates. The result follows from Lemma 3.1. ∎
Our proof of Theorem 4.5 relies on viewing the finite field $\mathbf{F}_{q^{n}}$ as a module over the modular group algebra of the Galois group $G=\operatorname{Gal}(\mathbf{F}_{q^{n}}/\mathbf{F}_{q})$, generated by the Frobenius endomorphism $\operatorname{Fr}$. More precisely, we consider the ring $\mathbf{F}_{q}[G]=\mathbf{F}_{q}[x]/(m(x))$, where $m(x)=x^{n}-1$ is the minimal and at the same time characteristic polynomial of $\operatorname{Fr}$ (see [LN83, proof of Theorem 2.35]). We can express the trace in terms of the polynomial
@@
since, for an element $\alpha\in\mathbf{F}_{q^{n}}$, we have $\operatorname{Tr}(\alpha)=\Phi(\operatorname{Fr})\alpha$. Consider now the module homomorphism
$\eta\colon\mathbf{F}_{q}[G]\to\mathbf{F}_{q^{n}},\quad t\mapsto t\cdot\alpha.$
Note that $\eta$ is an isomorphism precisely when $\alpha$ is a normal element in $\mathbf{F}_{q^{n}}$. Write $t=\sum_{i=0}^{n-1}t_{i}\operatorname{Fr}^{i}$ and let $\epsilon(t)=\sum_{i=0}^{n-1}t_{i}$ be the augmentation of $t$. Then
$\operatorname{Tr}(t\cdot\alpha)=\sum_{i=0}^{n-1}t_{i}\operatorname{Tr}(\operatorname{Fr}^{i}(\alpha))=\epsilon(t)\cdot\operatorname{Tr}(\alpha).$
For a normal $\alpha$, we have $\operatorname{Tr}(\alpha)\neq 0$, and so elements of trace 00 in $\mathbf{F}_{q^{n}}$ correspond precisely to the image of the augmentation ideal under $\eta$. This ideal is generated by $\operatorname{Fr}-1$, so traceless elements in $\mathbf{F}_{q^{n}}$ are generated by $(\operatorname{Fr}-1)(\alpha)=\alpha^{q}-\alpha$ as a module over $\mathbf{F}_{q}[G]$. The sharply traceless ones among these can be understood as follows.
Lemma 4.7.
Let $\alpha$ be a normal element in $\mathbf{F}_{q^{n}}$ and $\pi$ a polynomial in $\mathbf{F}_{q}[x]$. Then $\beta=\pi(\operatorname{Fr})(\operatorname{Fr}-1)(\alpha)$ is sharply traceless if and only if $\pi$ is coprime to $\Phi$.
@@
Suppose first that $\pi$ is coprime to $\Phi$. By Bézout, there are polynomials $a,b$ in $\mathbf{F}_{q}[x]$ so that $a\pi+b\Phi=1$. Evaluating in $\operatorname{Fr}$ and applying to $(\operatorname{Fr}-1)(\alpha)$, we obtain $a(\operatorname{Fr})(\beta)=(\operatorname{Fr}-1)(\alpha)$. Hence if some polynomial $c$ in $\mathbf{F}_{q}[x]$ satisfies $c(\operatorname{Fr})(\beta)=0$, then also $c(\operatorname{Fr})(\operatorname{Fr}-1)(\alpha)=0$, and since $\alpha$ is normal, it follows that $\Phi$ divides $c$. Thus $\beta$ is sharply traceless.
@@
Conversely, suppose that $\beta=\pi(\operatorname{Fr})(\operatorname{Fr}-1)(\alpha)$ is sharply traceless element. Note
$\frac{\Phi}{\gcd(\pi,\Phi)}(\operatorname{Fr})(\beta)=m(\operatorname{Fr})\frac{\pi}{\gcd(\pi,\Phi)}(\operatorname{Fr})(\alpha)=0.$
As $\beta$ is sharply traceless, it follows that $\gcd(\pi,\Phi)=1$. ∎
Note that two polynomials $\pi,\pi^{\prime}$ in $\mathbf{F}_{q}[x]$ determine the same sharply traceless element $\beta$ if and only if $(\operatorname{Fr}-1)(\pi-\pi^{\prime})(\operatorname{Fr})\alpha=0$, which is equivalent to saying that $\Phi$ divides $\pi-\pi^{\prime}$.
Corollary 4.8.
Sharply traceless elements in $\mathbf{F}_{q^{n}}$ are in bijective correspondence with the group of units of $\mathbf{F}_{q}[x]/(\Phi(x))$. [^3] The same argument gives that normal elements in $\mathbf{F}_{q^{n}}$ are in bijective correspondence with the group of units of $\mathbf{F}_{q}[x]/(m(x))$.
The number of these elements can be computed as follows. Let $\operatorname{Div}(\Phi)$ be the set of prime factors of $\Phi$. For a given $\pi\in\operatorname{Div}(\Phi)$, let $\operatorname{mult}(\pi)$ be the multiplicity of $\pi$ in $\Phi$. Then
$\frac{\mathbf{F}_{q}[x]}{(\Phi(x))}\cong\prod_{\pi\in\operatorname{Div}(\Phi)}\frac{\mathbf{F}_{q}[x]}{(\pi(x)^{\operatorname{mult}(\pi)})}.$
The invertible elements in each factor are polynomials that are coprime to $\pi$, so their number is
$q^{\deg(\pi)\operatorname{mult}(\pi)}-q^{\deg(\pi)(\operatorname{mult}(\pi)-1)}=q^{\deg(\pi)\operatorname{mult}(\pi)}(1-q^{-\deg(\pi)}).$
Note that $n-1=\deg(\Phi)=\sum_{\pi\in\operatorname{Div}(\Phi)}\deg(\pi)\operatorname{mult}(\pi)$, implying the following.